A 64-kg person riding on a roller coaster starts at rest and reaches a final speed of 26 m/s at the bottom of its first hill. What is the roller coaster’s height?
1
Expert's answer
2019-06-11T10:01:02-0400

The law of energy conservation:

"mgh=\frac{mv^2}{2}\\
=> h=\frac{v^2}{2g} = \frac{26^2}{2\cdot 9,81} = 34.45m"